A free atom of iron emits K base α X-rays of energy 6.4 keV. Calculate the recoil kinetic of the atom. Mass of an iron atom = 9.3 × 10^-20 kg.

Grade:10

1 Answers

Jitender Pal
9 years ago
Here energy of photon = E E = 6.4 KeV = 6.4 * 10^3 ev Momentum of Photon = E/C = 6.4 * 10^3/3 * 10^8 = 3.41 * 10^-24 m/sec. According to collision theory of momentum of photon = momentum of atom ∴ Momentum of Atom = P = 3.41 * 10^-24 m/sec ∴ Recoil K.E. of atom = P^2 /2m ⇒ (3.41 * 10^26)^2 eV/(2)(9.3 * 10^-26 * 1.6 * 10^-19) = 3.9 eV [1 joule = 1.6 * 10^-19 ev]